OpenStudy (anonymous):

solve √6x-1 =√5 <--(everything to the left of the = is under the radicall and the 5 is under a differnt radical)

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{6x-1}=\sqrt{5}\] \[\Large \left(\sqrt{6x-1}\right)^2=\left(\sqrt{5}\right)^2\] \[\Large 6x-1=5\] Can you take it from here?